Required information The following information applies to the questions displayed below.] Bethany incurred $29,750 in research and experimental costs for developing a specialized product during July of year 1 Bethany went through a lot of trouble and spent $26,750 in legal fees to receive a patent for the product in August of year 3. Bethany expects the patent to have a remaining useful life of 10 years. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.) . What amount of research and experimental expenses for year 1, year 2, and year 3 may Bethany deduct if she elects to amortize the xpenses over 60 months? Research and experimental expenses Year
Answer #1

eUAAAAASUVORK5CYII=

Workings

Monthly straight-line amortization =$29750/60=495.83

Year 1 straight-line amortization July through December =495.83*6=$2975.00

Year 2 straight-line amortization Jan through December =495.83*12=$5950

Bethany’s business months during year 3 before patent is issued in August =7 months

Year 3 straight-line amortization on 495.83*7=$3471